The square of a number is 10 more than 3 times the number.Find the number.Please Explain.
Answer by rothauserc(4718)

The square of a number is 10 more than 3 times the number.Find the number.Please Explain.
let x = the number then we have
x^2 = 3x + 10 and
x^2 -3x -10 = 0
factor the above equation
(x-5) * (x+2) = 0, then
(x-5)=0 and (x+2)=0
x can be 5 or -2